OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplifythe square root of 320.and provide the simplified form. @hartnn

OpenStudy (anonymous):

one way without sentences: \[\sqrt{320} = \sqrt{64 * 5} = \sqrt{64}*\sqrt{5} = 8\sqrt{5}\]

OpenStudy (anonymous):

@ash2326 is there another way?

OpenStudy (chmvijay):

 since 4 one is smaller u know the square root of 4 is 2 its easy to start with smaller integer :) hope it helps you